Shift Summary Day Tuesday 26/Oct/99

Herman B. White (hwhite@fnal.gov)
Tue, 26 Oct 1999 14:28:49 -0500 (CDT)


Shift Summary for Tuesday Day, 26/Oct/99

On Shift: H. White

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Shutdown

========================================================================
Accesses during shift: Tour, Interlock security test, High Bay Area
equipment review.

========================================================================
Overview of shift:

08:30 AM walkthrough.

No activity in the hall. A Heating & Air
conditioning service technician is working on the outside south air
conditioning unit, as well as the air handlers inside the electronics area
(after I let him enter).

09:30 AM
There was a meeting to review plans for removing materials off the
muon shielding blocks. This is required so that some shielding blocks
can be removed to extract the BA. Present: Koecher, Beverly,
Hsiung, Jensen, White). Decisions were made. If you have a concern, check
with Bud, Bob, or Doug.

10:00 AM
NA48 visitors, escorted by Hogan, access the hall for a tour.

10:30 AM Gas checklist complete.

11:00 AM Walkthrough and assisted operations crew with enclosure
interlock for security test. The duration of this test is not currently
known. It is expected to last an hour or so. Check with BobT or the main
control room for further information.

11:15 AM The NM3SP1A and NM3SP1B sump pumps are in alarm. There is water
in the pumping pit. The building mananger was notified, (and he visited).
He issued an investigate and repair work order for building services.
He also talked with the services people.
After the interlock test, a duty mechanic could access the enclosure. If
not today, then tomorrow morning.

Next, and subsequent walk throughs, take a look at the NM3 sump pumps, and
make sure no water is on the floor!

All other systems are quiet.

Wah awaits access to continue removing phototubes from HV bank.